### v2

# A data-driven framework for team selection in Fantasy Premier League

Abstract
Fantasy football is a billion-dollar industry with millions of participants. Under a fixed budget, managers select squads to maximize future Fantasy Premier League (FPL) points. This study formulates lineup selection as data-driven optimization and develops deterministic and robust mixed-integer linear programs that choose the starting eleven, bench, and captain under budget, formation, and club-quota constraints (maximum three players per club). The objective is parameterized by a hybrid scoring metric that combines realized FPL points with predictions from a linear regression model trained on match-performance features identified using exploratory data analysis techniques. The study benchmarks alternative objectives and cost estimators, including simple and recency-weighted averages, exponential smoothing, autoregressive integrated moving average (ARIMA), and Monte Carlo simulation. Experiments on the 2023/24 Premier League season show that ARIMA with a constrained budget and a rolling window yields the most consistent out-of-sample performance; weighted averages and Monte Carlo are also competitive. Robust variants improve some objectives but are not uniformly superior. The framework provides transparent decision support for fantasy roster construction and extends to FPL chips, multi-week rolling-horizon transfer planning, and week-by-week dynamic captaincy.
keywords:
data science and management , fantasy football , data-driven optimization , time series forecasting , machine learning
[…]
1 Introduction
Fantasy sports are a billion-dollar industry, with the market size expected to reach \$84.98 billion by 2032111https://www.skyquestt.com/report/fantasy-sports-market/. Fantasy football (soccer) is a leading segment: the official Fantasy Premier League (FPL) alone engaged over 11 million users in 2025222https://fantasy.premierleague.com/. In FPL, managers assemble a 15-player squad (two goalkeepers, five defenders, five midfielders, and three forwards) within a fixed budget of £100.0 m. Player prices evolve with performance and demand, and only the starting XI score in a given gameweek, with the captain’s points doubled; optional season-limited “chips” can alter scoring in specific weeks. Although the rules are simple, the weekly decision problem, which selects a cost-feasible starting XI and a captain under formation and club-quota constraints, with uncertain future performance, is nontrivial. With a global user base exceeding tens of millions and an industry valuation in the billions of dollars, the economic and social impact of fantasy sports cannot be overlooked. For participants, optimal decision-making translates into both competitive success and, in many cases, tangible financial rewards.
Against this backdrop, fantasy football has attracted attention across research and practitioner communities. Qualitative studies examine impacts on related markets and behavior, including merchandise sales [drayer2010effects], game attendance [nesbit2010impact], and participant well-being [wilkins2021exploring]. Quantitative research has explored performance prediction and strategy design using econometric and machine learning techniques. The increasing availability of public data and analytics tooling has further lowered barriers to empirical work and large-scale evaluation in this domain.
Despite this progress, most quantitative FPL studies emphasize forecasting weekly player points, with comparatively less attention to integrating those forecasts into a decision model that captures the full set of FPL constraints and jointly selects the starting XI and captain. Decision-analytic approaches exist [matthews2012competing], but captaincy is typically handled outside the optimization pipeline (e.g., via crowds or heuristics) rather than optimized jointly [bhatt2019should]. Moreover, robustness to estimation error is rarely addressed, and there is limited comparative evidence on how alternative objective functions, selection budget, or cost-vector estimation methods affect out-of-sample outcomes within a unified optimization framework.
Accordingly, we concentrate on two practical concerns: (i) converting noisy performance forecasts into an actionable weekly decision that jointly determines the starting XI and captain under budget, formation, and club-quota constraints; and (ii) mitigating sensitivity to estimation error and week-to-week variance. We address these by coupling integer programming models (with an explicit captaincy variable and operational constraints) with data-driven cost estimation and a hybrid scoring metric, and by evaluating a robust variant designed to hedge against misspecification. We also studied how budget constraints can change the results.
Generally, this paper makes the following contributions:
  • We formulate FPL lineup selection as a data-driven optimization problem and develop deterministic and robust integer models that choose the optimal starting XI, bench players, and the captain under budget, formation, and club-quota constraints.
  • We propose a hybrid scoring metric that combines realized FPL points with predictions from a ridge regression model trained on explainable match-performance features.
  • We build a unified pipeline that benchmarks alternative objectives and multiple forecasting approaches (including averaging techniques, exponential-smoothing, ARIMA, and simulation-based estimators) and provide an empirical assessment on the 2023/24 Premier League season with managerial insights into emergent formations and player selection patterns. We also implemented a rolling window approach and analyzed how constraining the budget for the starting XI affects the out-of-sample performance. To the best of our knowledge, this is the first study to provide a comprehensive analysis of data-driven team selection in FPL, leveraging a hybrid predictive metric within a robust decision-making framework.
  • All datasets, preprocessing scripts, and experiment notebooks are publicly released in a GitHub repository [^3] https://github.com/DanialRamezani/Data-Driven-FPL to ensure full reproducibility and to encourage further research and practical adoption.
The remainder of the paper is organized as follows. Sec. 2 reviews related work. Sec. 3 presents preliminaries and the optimization models. Sec. 4 introduces the proposed framework. Sec. 5 details the cost-vector estimation strategies and alternative objectives. Sec. 6 describes the dataset and feature construction. Sec. 7 shows exploratory data analysis results. Sec. 8 reports the empirical results and discusses the optimal lineups and formations. Sec. 9 concludes and outlines directions for future research.
2 Literature review
Several strands of prior work relate to our study [ati2024using]. First, decision-analytic and optimization approaches frame fantasy team management as an explicit choice problem. [matthews2012competing] model the FPL task as a belief-state Markov decision process with a tractable Bayesian Q-learning algorithm, casting weekly choices as a multi-dimensional knapsack and achieving around top-percentile performance against millions of human players. Extending optimization ideas to the NFL, [becker2016analytical] formulate a mixed-integer program to support draft selections and weekly lineup management using historical data and predictions. In FPL, [gupta2019time] propose a linear selection model for squad construction; while intentionally simple and omitting some operational constraints (e.g., optimal formation), they pair it with a hybrid RNN–ARIMA forecaster to supply objective inputs. More recently, metaheuristics and integrated pipelines have appeared: [aribo2024machine] apply a genetic algorithm for team selection and budget allocation, and [venter2024optimisation] combine optimization with forecasting in a practical system that attained a top-$4\%$ rank in the 2021/22 season. Taken together, these studies underscore the promise of explicit decision models, though they differ in how fully they encode FPL rules and whether captaincy is optimized jointly with the starting XI.
A second line of research develops predictive models for player performance, often treating the decision task separately. For NFL quarterbacks, [lutz2015fantasy] evaluated neural networks and support vector machines to improve score prediction, and [steenkiste2015finding] compared linear regression, random forests, and multivariate adaptive regression splines for next-week outcomes. Hybrid time-series ideas enter the literature with [gupta2019time], who fuse a recurrent neural network with ARIMA to estimate the cost vector used downstream. Within FPL, [bangdiwala2022using] assessed several machine-learning models and reported linear regression as the best performer on their dataset, while [pokharel2022fantasy] used an XGBoost regressor to study selection and transfer strategies through a return-on-investment lens and examine the effects of midweek cup fixtures. Deep sequence models have also been explored: [lombu2024predicting] compared convolutional neural networks with long short-term memory networks using form from the previous five matches, finding LSTM superior. Overall, the forecasting literature is methodologically rich, yet downstream optimization is often decoupled from prediction.
Complementing model-based prediction, researchers have examined crowd signals and ancillary tasks that influence managerial choices and market dynamics. [bhatt2019should] analyzed Twitter data for captain selection and found that the “wisdom of the crowd” outperforms expert analyst recommendations in their setting. In a related operational facet, [khamsan2019handling] addressed highly imbalanced classification when predicting virtual player price changes in FPL, where the majority class corresponds to no change.
Positioning this study within the literature, we bridge these strands by integrating data-driven forecasting with an integer optimization model that jointly selects the starting XI and the captain under full FPL constraints, and by assessing robustness to estimation error within the same framework. This unified treatment enables a comparative evaluation of objective functions and cost-vector estimators, linking predictive accuracy to operational decision quality.
3 Preliminaries and mathematical formulations
An FPL squad consists of 15 players: two goalkeepers (GK), five defenders (DEF), five midfielders (MID), and three forwards (FWD). Each gameweek, only the starting XI scores; one of the XI is named captain and receives double points each week. At most three players may come from any single Premier League team, and the total squad value cannot exceed £100.
For simplicity, gamers who select players will be referred to as team owners, while professional Premier League footballers are called players. As previously mentioned, the goal of this paper is to select players in a way that maximizes the total points in the upcoming weeks. Therefore, special and one-time cards are not taken into account. The selection of the starting eleven can be formulated as the following integer programming model:
  • $n$: total number of players.
  • $c_{j}$: expected points for player $j$.
  • $b$: total budget for the starting eleven.
  • $v_{j}$: value (price) of player $j$.
  • $\text{min\_limit}_{k}$: minimum number of players to be selected from position $k$.
  • $\text{max\_limit}_{k}$: maximum number of players to be selected from position $k$.
  • $\text{Team}_{t}$: set of players from team $t$.
Decision variables:
  • $x_{j}\in\{0,1\}$ for $j=1,2,\dots,n$: $x_{j}=1$ if player $j$ is selected in the starting eleven; 0 otherwise.
  • $y_{j}\in\{0,1\}$ for $j=1,2,\dots,n$: $y_{j}=1$ if player $j$ is selected as captain; 0 otherwise.
Objective:
$\max\;Z=\sum_{j=1}^{n}c_{j}x_{j}+\sum_{j=1}^{n}c_{j}y_{j}$ (1)
The objective maximizes the expected reward of the team. In FPL, the captain’s score is doubled; modeling captaincy with $y_{j}$ adds a second $c_{j}$ term when player $j$ is captain.
Subject to:
$\sum_{j=1}^{n}x_{j}=11$ (2)
Eq. 2 indicates that exactly eleven starters are selected.
$\sum_{j=1}^{n}v_{j}x_{j}\leq b$ (3)
Eq. 3 indicates that the total cost of the starting eleven does not exceed the budget $b$.
$\sum_{j=1}^{n}y_{j}=1$ (4)
Eq. 4 indicates that exactly one captain is chosen.
$y_{j}\leq x_{j},\quad\forall j=1,2,\dots,n$ (5)
Eq. 5 indicates that the captain must be among the selected starters.
$\sum_{j\in\text{Position}_{k}}x_{j}\geq\text{min\_limit}_{k},\quad\forall k\in\{\text{GK},\text{DEF},\text{MID},\text{FWD}\}$ (6)
Eq. 6 enforces the minimum positional counts for a legal formation.
$\sum_{j\in\text{Position}_{k}}x_{j}\leq\text{max\_limit}_{k},\quad\forall k\in\{\text{GK},\text{DEF},\text{MID},\text{FWD}\}$ (7)
Eq. 7 enforces the maximum positional counts for a legal formation.
$\sum_{j\in\text{Team}_{t}}x_{j}\leq 3,\quad\forall t=1,2,\dots,20$ (8)
Eq. 8 ensures that at most three players may be selected from any single Premier League club.
In this work, various methods are analyzed to estimate the best choice of $c_{j}$ for out-of-sample performance. Although decision-makers are free to define any budget below 100, in this study, we set $b=83.5$. The rationale is straightforward: the model must also account for four reserve players, and in FPL the cheapest players typically cost around four million pounds each, totaling 16 million. An additional 0.5 million is added as a buffer to prevent budget violations. Thus, the formulation focuses on selecting the best fixed starting squad, while team owners are free to choose four inexpensive reserve players for a total of up to £ 16.5 million.
3.1 Optimizing bench
In the official FPL, managers register 15 players with positional totals $2\text{--}5\text{--}5\text{--}3$ (two goalkeepers, five defenders, five midfielders, three forwards). We therefore optimize the four bench slots subject to budget, team quota, and positional complement constraints so that the combined roster (starters + bench) satisfies the official totals.
Additional decision variables:
  • $x^{b}_{j}\in\{0,1\}$ for $j=1,2,\dots,n$: $x^{b}_{j}=1$ if player $j$ is selected for the bench; 0 otherwise.
Additional parameters:
  • $b^{\text{bench}}$: budget available for the bench (100b100-b).
  • $r_{k}$: required number of players in position $k$ for the bench.
  • $s_{t}$: number of players from team $t$ already selected in the starting XI.
  • $S$: set of players already selected in the starting XI.
Figure 1: The proposed framework. Modeling produces expected points, which feed the optimization to select XI and captain under FPL constraints.
Bench objective:
$\max\;Z^{bench}=\sum_{j=1}^{n}c_{j}x^{b}_{j}$ (9)
Eq. 9 maximizes the expected contribution of bench players, providing insurance when a starter does not feature.
Bench constraints:
$\sum_{j\in\text{Position}_{k}}x^{b}_{j}=r_{k},\quad\forall k\in\{\text{GK},\text{DEF},\text{MID},\text{FWD}\}$ (10)
Eq. 10 enforces positional requirements for the bench so the full 15-man roster satisfies FPL’s 2–5–5–3 totals.
$\sum_{j=1}^{n}v_{j}x^{b}_{j}\leq b^{\text{bench}}$ (11)
Eq. 11 keeps the bench within the remaining budget $b^{\text{bench}}=100-b$.
$\sum_{j\in\text{Team}_{t}}x^{b}_{j}+s_{t}\leq 3,\quad\forall t=1,2,\dots,20$ (12)
Eq. 12 extends the three-player-per-team restriction to the full squad (starters + bench).
$x^{b}_{j}=0,\quad\forall j\in S$ (13)
Eq. 13 prevents duplication: no player may appear in both the starting XI and the bench.
3.2 Robust optimization
Robust optimization is a decision-making framework under uncertainty. Unlike stochastic programming, which requires perfect knowledge of the distribution of the uncertain parameters and scenario generation, robust optimization focuses on ensuring that the model solution remains feasible and performs well under the worst-case realization of uncertainty.
Let the actual expected points be uncertain and lie in a box uncertainty set:
$\mathcal{U}=\left\{c\in\mathbb{R}^{n}:c_{j}\in[\bar{c}_{j}-d_{j},\bar{c}_{j}+d_{j}]\;\;\forall j\right\},$ (14)
In Eq. 14, $\bar{c}_{j}$ is the nominal expected score and $d_{j}\geq 0$ is the uncertainty margin for player $j$.
Finally, the worst-case objective is formulated as follows:
$\max_{x,y}\;\min_{c\in\mathcal{U}}\sum_{j=1}^{n}c_{j}(x_{j}+y_{j}).$ (15)
Eq. 15 ensures that the selected team maximizes the worst-case total score, making the solution robust against possible overestimation in the expected scores and week-to-week variance.
4 The proposed framework
Figure 1 shows the proposed end-to-end pipeline from the merged FPL dataset to a team recommendation. We begin with a player–gameweek (GW) panel, verify types, remove duplicates, and guard against leakage by excluding any same-GW outcomes (e.g., bonus/BPS) from predictors. We fix a train→test split (GW1–26 → GW27–38) and treat missing player–GW rows as unavailability rather than an explicit injury flag. For the target GW we also built a compact table schema for the optimizer—name, position, team, value, and a slot for expected points—with the slot filled by the models later; modeling inputs are standardized.
[…]
The modeling stage fits five families: (1) averages/smoothing, (2) ARIMA variants, (3) linear trend regression, (4) simulation (bootstrap and parametric), and (5) the hybrid that combines recent outcomes with the Ridge feature model. These produce per-player expected points $c_{j}$ (and, when applicable, uncertainty widths $d_{j}$).
Finally, a 0–1 integer program selects the starting XI and captain under FPL rules—budget, valid formations, and at most three players per club—with the captain required to start. A bench is then chosen from the remaining budget to complete a 2–5–5–3 squad under the same club quota. Besides maximizing expected points, we also report surrogate XI objectives (xGI-xGC or ICT) and an optional robust variant that penalizes scores by uncertainty ($c_{j}\pm d_{j}$). We output the recommended XI, captain, and bench, along with formation trends, weekly and cumulative points, method leaderboards, and budget/rolling-window sensitivity analyses.
5 Estimating cost vectors
This section details how we estimate the per–player expected points $c_{j}$ used by the optimization models in Sec. 3. All methods consume historical weekly points up to a fixed split week and return a single scalar summary per player. Let the season length be $N=38$ gameweeks and the train/test split occur at gameweek $\tau$ (train: weeks $1{:}\tau$; test: weeks $\tau{+}1{:}N$). For a player with time series $\{p_{1},\dots,p_{\tau}\}$, the goal is to produce a representative value $c$ summarizing expected future performance.
5.1 Averaging-based estimators
We begin with simple summaries of $\{p_{t}\}$.
Simple average
$\widehat{c}^{\,\text{SA}}\;=\;\frac{1}{\tau}\sum_{t=1}^{\tau}p_{t}$ (16)
Eq. 16 treats all observed weeks equally.
Weighted average (recency emphasis)
Let the linear weights be $w_{t}=t\big/\sum_{i=1}^{\tau}i$.
$\widehat{c}^{\,\text{WA}}\;=\;\sum_{t=1}^{\tau}w_{t}\,p_{t}$ (17)
Eq. 17 indicates that more recent weeks receive larger weights; week $\tau$ has the highest weight.
Exponential smoothing (Holt’s linear trend)
We fit a level–trend model (no seasonality):
$\hat{p}_{t}$ $=\ell_{t-1}+b_{t-1}$ (18)
$\ell_{t}$ $=\alpha\,p_{t}+(1-\alpha)(\ell_{t-1}+b_{t-1})$ (19)
$b_{t}$ $=\beta\,(\ell_{t}-\ell_{t-1})+(1-\beta)b_{t-1}$ (20)
In the above equations, $\ell_{t}$ and $b_{t}$ denote level and trend; $\alpha,\beta\in[0,1]$ are estimated from data. Forecast the remaining weeks $\{\hat{p}_{\tau+1},\dots,\hat{p}_{N}\}$ and summarize by
$\widehat{c}^{\,\text{ES}}\;=\;\frac{1}{N-\tau}\sum_{k=1}^{N-\tau}\hat{p}_{\tau+k}.$ (21)
Eq. 21 averages the future forecasts to produce a single score.
5.2 Simulation-based estimators
We approximate future outcomes by sampling from historical behavior.
Bootstrap (nonparametric)
Draw $B$ resamples of length $N-\tau$ with replacement from $\{p_{1},\dots,p_{\tau}\}$; average within resamples, then across resamples:
$\widehat{c}^{\,\text{BOOT}}\;=\;\frac{1}{B}\sum_{b=1}^{B}\left(\frac{1}{N-\tau}\sum_{k=1}^{N-\tau}p^{*(b)}_{k}\right)$ (22)
Eq. 22 makes no distributional assumptions; $p^{*(b)}_{k}$ are bootstrap draws.
Monte Carlo (parametric)
Assume $p_{t}$ are i.i.d. with mean $\bar{p}=\frac{1}{\tau}\sum_{t=1}^{\tau}p_{t}$ and variance $s^{2}$. Simulate $B$ future paths from a chosen family (e.g., Gaussian or truncated Gaussian) with parameters $(\bar{p},s^{2})$ and compute the same summary as in (22) to obtain $\widehat{c}^{\,\text{MC}}$.
5.3 Autoregressive integrated moving average (ARIMA)
We fit an $\text{ARIMA}(p,d,q)$ model to $\{p_{t}\}_{t=1}^{\tau}$:
$\phi(B)\,(1-B)^{d}\,p_{t}\;=\;\theta(B)\,\varepsilon_{t}$ (23)
In Eq. 23, $B$ is the backshift operator; $\phi(\cdot)$ and $\theta(\cdot)$ are AR and MA polynomials; $\varepsilon_{t}$ is white noise. Rolling forecasts $\{\hat{p}_{\tau+1}$, $\dots$, $\hat{p}_{N}\}$ are summarized by
$\widehat{c}^{\,\text{ARIMA}}\;=\;\frac{1}{N-\tau}\sum_{k=1}^{N-\tau}\hat{p}_{\tau+k}.$ (24)
Eq. 24 uses the average of multi-step forecasts as a single cost value.
5.4 Linear regression (time-on-trend)
Regress points on week index $w_{t}=t$ for $t=1,\dots,\tau$:
$p_{t}\;=\;\beta_{0}+\beta_{1}w_{t}+\varepsilon_{t}$ (25)
In Eq. 25, $\beta_{0},\beta_{1}$ are OLS estimates; $\varepsilon_{t}$ are residuals. Predict weeks $w_{\tau+1},\dots,w_{N}$ and summarize:
$\widehat{c}^{\,\text{LR}}\;=\;\frac{1}{N-\tau}\sum_{k=\tau+1}^{N}\hat{p}_{k}.$ (26)
Eq. 26 averages the linear-trend forecasts to a single score.
5.5 Hybrid method (ridge + realized points)
The hybrid approach augments weekly points with match-performance features that are not directly scored by FPL but correlate with future returns (e.g., ICT components, starts, shots). Let $\mathbf{X}$ be feature vectors and $\mathbf{y}$ be targets (weekly points) up to week $\tau$. Standardize features and fit ridge regression:
$\hat{\mathbf{w}}\;=\;\arg\min_{\mathbf{w}}\left\{\left\|\mathbf{y}-\mathbf{Xw}\right\|_{2}^{2}+\alpha\left\|\mathbf{w}\right\|_{2}^{2}\right\}$ (27)
In Eq. 27, $\alpha\geq 0$ controls shrinkage. Use the fitted model to predict per–player season-forward means $\widehat{y}$ (e.g., by averaging predicted weeks $\tau{+}1{:}N$). Combine normalized realized and predicted values:
$\widehat{c}^{\,\text{HYB}}\;=\;(1-\lambda)\,y^{\text{norm}}\;+\;\lambda\,\widehat{y}^{\,\text{norm}},\quad\lambda\in[0,1].$ (28)
In Eq. 28, $y^{\text{norm}}$ and $\widehat{y}^{\text{norm}}$ are per–player normalizations (e.g., z scores or min–max). We evaluate $\lambda=\tfrac{1}{3}$ (2:1 favoring realized) and $\lambda=\tfrac{2}{3}$ (1:2 favoring predicted).
5.6 Alternative objective surrogates
For completeness, we also replace points with proxy targets.
ICT index maximization
$\max\sum_{j=1}^{n}\text{ICT}_{j}\,(x_{j}+y_{j})$ (29)
In Eq. 29, $\text{ICT}_{j}$ denotes the expected ICT score of player $j$; the captain contribution is doubled via $y_{j}$.
Attack–defense trade-off (EGI-EGC)
$\max\sum_{j=1}^{n}\big(\text{EGI}_{j}-\text{EGC}_{j}\big)\,(x_{j}+y_{j})$ (30)
In Eq. 30, $\text{EGI}_{j}$ is expected goal involvement; $\text{EGC}_{j}$ is expected goals conceded for player $j$. Higher values prefer attacking contributors while discouraging players expected to concede.
6 Dataset and experimental setup
We use the publicly available FPL dataset [^4] https://github.com/vaastav/Fantasy-Premier-League/, which provides a player–gameweek panel for each Premier League season (38 gameweeks). FPL awards points according to role-specific rules (e.g., clean sheets for defenders/goalkeepers; goals/assists for attackers), while some events apply to all positions (e.g., minutes played, cards) [^5] https://www.premierleague.com/news/2174909/. Feature names follow the official FPL statistics glossary [^6] https://fantasy.premierleague.com/statistics/.
Figure 2: Correlation between selected features and total points by positions.
Figure 3: Pearson correlation heatmap of modeling features.
Figure 4: RainCloud plots of top numeric features by position.
(a) Goalkeepers
(b) Defenders
(c) Midfielders
(d) Forwards
Figure 5: SHAP value summaries for key features by position.
The unit of analysis is a player–gameweek. Unless stated otherwise, we use the 2023/24 season with $N=38$ gameweeks and a fixed split at gameweek $\tau$: training on weeks $1{:}\tau$ and evaluating on weeks $\tau{+}1{:}N$ (consistent with Sec. 5). Player identity is the FPL unique player code; club membership is the contemporaneous club in that gameweek (relevant for the club–quota constraints in Sec. 3). Prices are scaled to millions, as in our code (value divided by 10), and are used downstream for the budget constraint.
Our outcome is gameweek total points, denoted $p_{t}$ in Sec. 5. To avoid target leakage, we exclude contemporaneous outcomes and referee events as predictors for all models: goals scored, clean sheets, goals conceded, penalties missed, saves, penalties saved, own goals, yellow cards, red cards, assists, team scores (team h score, team a score), and the bonus system (bps, bonus). We also drop schedule/row identifiers that do not encode player features (opponent team, GW/round, kickoff time, fixture, was home) as modeling features. The target–week minutes is never used as a predictor. The resulting filtered table is used to construct the covariates for the hybrid/surrogate experiments below, while prices (values), teams, and positions are retained for the optimizer.
The following shows the detailed setup of the prediction models.
  1. (a)
    History-only forecasters include simple and weighted averages, exponential smoothing, ARIMA, linear trend, and simulation (bootstrap/normal Monte Carlo). These methods estimate each player’s expected points using only $\{p_{1},\ldots,p_{\tau}\}$; no exogenous features are consumed. We also retain the empirical standard deviation of past $p_{t}$ as a simple uncertainty proxy.
  2. (b)
    We form per–player aggregates over the training window for non–scoring, pre–deadline proxies that survive filtering: ict index,expected goals (xG), expected assists (xA), expected goal involvements (xGI), expected goals conceded (xGC), selected, starts.
    • For each position, we fit a ridge regression of
      total points on the proxies above (standardized), and combine the ridge prediction with the normalized historical mean of total points to obtain a hybrid score.
    • For interpretability, we also fit an XGBoost regressor on the same proxies and use SHAP to visualize influential features. This step is diagnostic only and is not fed to the optimizer in the reported runs.
    • Surrogates: maximize $\mathrm{EGI}-\mathrm{xGC}$ and use a robust ICT score $\emph{ict index}-\mathrm{sd}(\emph{ict index})$.
Unless stated otherwise, the integer program in Sec. 3 uses value for the budget, position for formation limits, and team for the $\leq 3$ per–club rule. The starting XI objective is chosen from the set above (expected points, hybrid score, EGI-xGC, robust ICT). The bench is ordered by the history-only expected points.
7 Exploratory data analysis
Fig. 2 shows Pearson correlations between the modeling features and gameweek total points by position, computed on the training window ($\mathrm{GW}\leq\tau$). Across roles, ict index is the strongest single proxy, and exposure signals such as starts (and selected) also correlate positively with points. For attackers (Forward, Midfielder), expected goal involvements and expected goals exhibit sizeable positive correlations; these signals are weaker but still positive for Defenders. For the Goal Keeper panel, starts and ict index dominate, with other proxies contributing modestly.
Fig. 3 highlights a clear “attacking” cluster. Features including xG, xA, xGI, and ict index are all strongly and mutually correlated, and each shows a solid positive relationship with total points. Starts and selected also track positively with points (proxies for minutes/availability), while value co-moves with performance and popularity. By contrast, xGC has the weakest—and likely slightly negative—association with total points and only mild links to the attacking features. Overall, the tight correlations among xG/xA/xGI/ICT indicate notable multicollinearity, so using regularization or a reduced set (e.g., keep xGI) would avoid redundancy.
Fig. 4 shows zero-inflation and heavy right tails for attacking metrics (xA, xG, xGI, ICT), position-dependent heteroscedasticity (GK/DEF tight; MID/FWD wide), and strong collinearity among attacking features, with value rising along the same gradient. This shape favors recency-aware, shrinkage forecasts with the rolling ARIMA and weighted averages because they 1) stabilize means in the presence of outliers/hauls, 2) adapt to non-stationary form and changing minutes, and 3) avoid exploding variance that simple means or Monte-Carlo draws suffer in heavy-tailed data. Collinearity among xG/xA/xGI/ICT motivates feature parsimony or ridge in the hybrid metric (often collapsing to xGI + regularization). On the optimization side, the distributions justify a budget- and formation-constrained IP: it allocates spend and captaincy toward the higher-mean, higher-upside MID/FWD slots while using cheaper, lower-variance GK/DEF to meet minima; an objective that rewards xGI and/or penalizes xGC aligns with the positional patterns seen.
Fig. 5 shows position-wise SHAP summaries from the diagnostic XGBoost: the horizontal spread indicates importance (wider = larger impact), and color encodes the raw feature value (pink = high, blue = low). It can be observed that the ICT index is the most influential predictor across DEF/MID/FWD, with higher ICT consistently pushing predictions up. For attackers, xGI (and its components xG, xA) strongly and positively drive the model, while starts mainly captures availability effects. On the defensive side, expected goals conceded (xGC) has a clear negative contribution (higher xGC leads to a lower output), especially for defenders, and starts is most important for goalkeepers. Selected% and value add comparatively little marginal signal once the other features are included.
8 Experimental results and discussion
(a) Averaging methods and robust counterparts
(b) Simulation methods (bootstrap vs. Monte Carlo)
Figure 6: Out-of-sample performance: averaging and simulation families.
Figure 7: Out-of-sample performance of ARIMA variants. Simpler models (ARIMA (0,0,1) and ARIMA (1,0,1)) outperform higher-order configurations.
Figure 8: Alternative objectives: ICT, robust ICT, and involvement (EGI-EGC).
(a) Hybrid over averaging methods
(b) Hybrid over simulation methods
Figure 9: Effect of the hybrid metric on averaging and simulation families.
(a) Hybrid vs. ARIMA (0,0,1)
(b) Hybrid ICT vs. ICT and linear regression
Figure 10: Hybridization applied to ARIMA and ICT/linear regression.
(a) Simple average
(b) Weighted average
Figure 11: Budget sensitivity: averaging methods (static teams).
(a) ARIMA (1,0,0) (rolling)
(b) ARIMA (1,0,1) (rolling)
Figure 12: Rolling selection: ARIMA (1,0,0) and ARIMA (1,0,1).
We evaluate models by out-of-sample cumulative points. Training uses gameweeks 1–26 of the 2023/24 season; evaluation spans gameweeks 27–38. In gameweek 29, many player records are missing; instead of discarding teams, we follow the previously explained bench substitution policy. All plots report cumulative team points (no chips). To avoid clutter, we show the best-performing variant within each method family.
Fig. 6 compares averaging-based estimators, and their robust counterparts, to simulation approaches (bootstrap and parametric Monte Carlo, each with B=1000B{=}1000 paths). It can be observed that (i) Weighted average (recency) consistently dominates simple averaging and exponential smoothing, indicating that recent form provides a useful bias signal for FPL points. (ii) Robustification does not help the averaging family: conservatism against point overestimation sacrifices upside without materially reducing downside weeks, so robust curves largely track or underperform their deterministic counterparts. (iii) Bootstrap vs. Monte Carlo are near indistinguishable across the horizon, suggesting that, at this granularity, distributional assumptions add little beyond the empirical resampling of historical points.
Fig. 7 summarizes nine ARIMA specifications. It can be observed that simpler ARIMA structures generalize better: ARIMA (0,0,1) and ARIMA (1,0,1) exceed 600 cumulative points, ARIMA (1,0,0) trails slightly, and performance degrades for more complex models (e.g., ARIMA (2,1,2)). This is consistent with the short, noisy weekly horizon where over-parameterized models overfit in-sample fluctuations that do not persist.
Fig. 8 contrasts ICT-based objectives with an attack-defense surrogate that maximizes expected goal involvement minus expected goals conceded (EGI-EGC). It can be observed that robust ICT modestly improves over deterministic ICT, unlike the averaging family, but the gain is small. The involvement proxy underperforms, likely because conceding risk is not individualized well at the player level and is already indirectly captured by team quality and minutes.
(a) ARIMA (0,0,1)
(b) ARIMA (1,0,0)
(c) ARIMA (1,0,1)
Figure 13: Budget sensitivity: ARIMA family (static teams).
(a) ICT
(b) Hybrid ICT (1:2)
(c) Monte Carlo
Figure 14: Budget sensitivity: ICT objectives and Monte Carlo (static teams).
(a) Simple average (rolling)
(b) Weighted average (rolling)
(c) ARIMA (0,0,1) (rolling)
Figure 15: Rolling selection: averaging family and ARIMA (0,0,1).
Figure 16: Best-performing models across families (cumulative points).
Figure 17: Comparing Results with [santorum2025FPL].
We examine the hybrid metric (Eq. (28)) applied to averaging/simulation families (Fig. 9) and to ARIMA and ICT (Fig. 10). Across families, the 1:2 setting (favoring predictions) is more reliable than 2:1 (favoring realized points): it typically provides a small uplift or remains competitive, whereas 2:1 often dilutes the signal by over-weighting historical noise. For ARIMA (0,0,1), hybridization does not improve the season-long trajectory, implying that the ARIMA residual structure already captures the predictable component. For ICT, the hybrid (1:2) variant offers a clear improvement over plain ICT, consistent with the hybrid’s ability to fuse direct non-scoring proxies with recent realized outcomes.
We next vary the starting-XI budget (bench receives the residual to reach £100m) with fixed teams (no weekly repick). Figures 1114 summarize trends across families. For simple averages, tighter starting budgets (down to $\sim£65$-£70m) can outperform the default £83.5m, consistent with stronger benches covering blanks and injuries. For weighted averages, reducing the starting budget generally hurts, suggesting the best performers (recent form) reside among premium or mid-premium assets. In the ARIMA family, ARIMA (0,0,1) is relatively insensitive across £75–£83.5m, while ARIMA (1,0,0) benefits from tighter starting budgets (£70–£80m); tightening the budget leads to a decline in the performance of the ARIMA(1,0,1) model. While the hybrid ICT model does not change much with varying budgets, the ICT score model achieves its highest performance with the default budget; the Monte Carlo model also benefits from a tighter budget, except at £60 million, where talent scarcity dominates.
We now re-pick the starting XI from a fixed 15-man roster each gameweek (no transfers), promoting bench players when their expected points exceed a starter’s and the full roster remains 22-55-55-33 compliant. A rolling policy helps the simple average (notably around £70-£80m) by enabling timely bench promotions; it has little effect on weighted averages and ARIMA (0,0,1). This, however, was not the case in ARIMA (1,0,1), where the rolling setting with £70-£80m outperformed the default model. Figures 12-15 summarize trends across families. For ARIMA (1,0,0), rolling selection is a major uplift, reaching almost $\sim$700 cumulative points with £70–£75m starting budgets, the strongest result in our study. This indicates that a modest trend model benefits disproportionately from active lineup refresh within a fixed squad.
Figure 16 reports the best-performing method within each model family. Table  LABEL:teb:summary provides the final cumulative values for each model in the above-mentioned figure and their comparative rankings. Figure 17 then benchmarks two models from this study, an ARIMA(1,0,0) forecaster with a rolling window (starting-XI budget of £70m) and a weighted-average forecaster, against the four approaches in [santorum2025FPL]. In brief, the Greedy algorithm builds a 15-player squad by ranking a cost-performance (“goodness”) index based on predicted mean season points, with the weekly XI chosen at random; Mean-MIP-random uses Mixed Integer Programming (MIP) on predicted mean scores to select the initial 15, again with a random weekly XI; Mean-MIP-rank keeps the MIP-optimized squad but selects the XI each Gameweek via MIP using adjusted, GW-specific scores (form, home/away, opponent difficulty); and Mean-MIP-online updates the predictions every Gameweek and re-optimizes the XI via MIP while holding the initial squad fixed.
Table 1: Final Cumulative Points and Ranks by Method
Method Final Total Rank
ARIMA (1,0,0) Rolling (Budget = 70) 704 1
Weighted Average 635 2
Hybrid Simple Avg. 1:2 (Higher Total Points) 561 3
Monte Carlo Simulation 545 4
Table 2: Weighted Average Team for Gameweek 27.
Name Team Position Value
Jordan Pickford Everton GK 4.6
Gabriel dos Santos Magalhães Arsenal DEF 5.2
William Saliba Arsenal DEF 5.7
Virgil van Dijk Liverpool DEF 6.4
Phil Foden Man City MID 8.1
Douglas Luiz Soares de Paulo Aston Villa MID 5.5
Bukayo Saka (c) Arsenal MID 9.1
Cole Palmer Chelsea MID 5.8
Mohamed Salah Liverpool MID 13.0
Ollie Watkins Aston Villa FWD 8.8
Dominic Solanke Bournemouth FWD 7.0
Matheus Cunha Wolves FWD 5.6
Jarrad Branthwaite Everton DEF 4.2
Ederson Santana de Moraes Man City GK 5.5
Conor Bradley Liverpool DEF 4.1
Table 3: ARIMA (1,0,0) Rolling (Budget = 70) Team for Gameweek 27
Name Team Position Value
Đorđe Petrović Chelsea GK 4.5
Gabriel dos Santos Magalhães Arsenal DEF 5.2
William Saliba Arsenal DEF 5.7
Chris Richards Crystal Palace DEF 3.9
Phil Foden Man City MID 8.1
Douglas Luiz Soares de Paulo Aston Villa MID 5.5
Anthony Gordon Newcastle MID 6.1
Bukayo Saka (c) Arsenal MID 9.1
Cole Palmer Chelsea MID 5.8
Ollie Watkins Aston Villa FWD 8.8
Dominic Solanke Bournemouth FWD 7.0
Erling Haaland Man City FWD 14.4
Jordan Pickford Everton GK 4.6
Pedro Porro Spurs DEF 5.8
Emerson Palmieri dos Santos West Ham DEF 4.4
Table 4: Monte Carlo Team for Gameweek 27.
Name Team Position Value
Đorđe Petrović Chelsea GK 4.5
Gabriel dos Santos Magalhães Arsenal DEF 5.2
William Saliba Arsenal DEF 5.7
Kieran Trippier Newcastle DEF 6.9
Phil Foden Man City MID 8.1
Douglas Luiz Soares de Paulo Aston Villa MID 5.5
Son Heung-min Spurs MID 9.6
Bukayo Saka (c) Arsenal MID 9.1
Mohamed Salah Liverpool MID 13.0
Ollie Watkins Aston Villa FWD 8.8
Dominic Solanke Bournemouth FWD 7.0
Alphonse Areola West Ham GK 4.2
Conor Bradley Liverpool DEF 4.1
Cameron Archer Sheffield Utd FWD 4.3
Amari’i Bell Luton DEF 3.9
Table 5: Hybrid Simple Average (1:2) Team for Gameweek 27.
Name Team Position Value
Alphonse Areola West Ham GK 4.2
William Saliba Arsenal DEF 5.7
Trent Alexander-Arnold Liverpool DEF 8.6
Kieran Trippier Newcastle DEF 6.9
Phil Foden Man City MID 8.1
Douglas Luiz Soares de Paulo Aston Villa MID 5.5
Pascal Groß Brighton MID 6.5
Jarrod Bowen West Ham MID 7.7
Cole Palmer Chelsea MID 5.8
Erling Haaland (c) Man City FWD 14.4
Ollie Watkins Aston Villa FWD 8.8
Jarrad Branthwaite Everton DEF 4.2
Conor Bradley Liverpool DEF 4.1
Cameron Archer Sheffield Utd FWD 4.3
Matt Turner Nott’m Forest GK 3.9
Table 6: Linear Regression Team for Gameweek 27.
Name Team Position Value
Ivo Grbić Sheffield Utd GK 4.5
Gabriel dos Santos Magalhães Arsenal DEF 5.2
Daniel Muñoz (c) Crystal Palace DEF 4.5
Conor Bradley Liverpool DEF 4.1
Phil Foden Man City MID 8.1
Bukayo Saka Arsenal MID 9.1
Cole Palmer Chelsea MID 5.8
Richarlison de Andrade Spurs MID 7.2
Ollie Watkins Aston Villa FWD 8.8
Rasmus Højlund Man Utd FWD 7.2
Jayden Danns Liverpool FWD 4.5
Pablo Sarabia Wolves MID 4.7
Jarrad Branthwaite Everton DEF 4.2
Chris Richards Crystal Palace DEF 3.9
Caoimhin Kelleher Liverpool GK 3.7
Table 7: ICT Score Team for Gameweek 27.
Name Team Position Value
James Trafford Burnley GK 4.5
Pedro Porro Spurs DEF 5.8
Kieran Trippier Newcastle DEF 6.9
Alfie Doughty Luton DEF 4.6
Phil Foden Man City MID 8.1
Pascal Groß Brighton MID 6.5
Bruno Borges Fernandes Man Utd MID 8.2
Bukayo Saka (c) Arsenal MID 9.1
Martin Ødegaard Arsenal MID 8.4
Ollie Watkins Aston Villa FWD 8.8
Dominic Solanke Bournemouth FWD 7.0
Alphonse Areola West Ham GK 4.2
Conor Bradley Liverpool DEF 4.1
Cameron Archer Sheffield Utd FWD 4.3
Amari’i Bell Luton DEF 3.9
As shown in Figure 17, both our ARIMA(1,0,0)-Rolling and weighted-average models outperform the baselines across most of the season. The panel also plots week-by-week differences (ours minus the best competing method); the negative dip in Gameweek 29 stems from data availability rather than model behavior. The only exceptions are Gameweeks 29 and 38, where the MIP-online approach is ahead; otherwise, our two models lead throughout.
Tables 27 present the teams selected by our top model families for Gameweek 27, revealing clear structural regularities across methodologies. Midfield-centric formations, particularly 3-5-2, dominate in almost all approaches, underscoring the models’ collective preference for midfielders’ attacking output and bonus potential under the FPL scoring system. The linear regression model leans toward a 3-4-3 configuration.
Regarding captaincy, Bukayo Saka is the armband choice for five of the six models, highlighting his centrality in the data-driven selection process. Cross-model consistency is also evident in the frequent inclusion of Phil Foden, Cole Palmer, and Ollie Watkins—the latter appearing in every best XI, consistent with Premier League’s Gameweek 27 review (“GW27 stats: Brave FPL managers rewarded for captaining Watkins”). [^7] https://www.premierleague.com/news/3917209 Arsenal assets, especially defenders Gabriel dos Santos Magalhães and William Saliba, recur prominently, reflecting both strong defensive metrics and attacking involvement.
In terms of premium allocation, Mohamed Salah and Erling Haaland are included across several lineups but seldom together, as their combined cost exceeds typical budget constraints unless projected returns justify the expenditure. The goalkeeper position remains the most variable: the ARIMA and Monte Carlo models favor Đorđe Petrović, the weighted average opts for Jordan Pickford, and the hybrid approach selects Alphonse Areola, illustrating contrasting valuation strategies between clean-sheet reliability and save-point potential.
Bench selections are largely consistent, with Conor Bradley and Cameron Archer repeatedly chosen as cost-efficient enablers. Overall, the model outputs exhibit a pronounced convergence toward big teams’ assets(e.g., Arsenal and Liverpool) and Aston Villa, balanced by strategic differentials in defense and goalkeeping designed to optimize both expected value and formation flexibility.
9 Conclusion
This paper presents an integer–programming framework with a robust variant that turns player-level forecasts into valid FPL lineups. The optimizer selects the starting XI, bench, and captain under budget, formation, and club limits, using data–derived expected points as inputs. To supply these inputs, we compared averaging methods, lightweight time–series forecasting, simulation techniques, and a hybrid approach that blends recent outcomes with feature-based predictions. We also examined the impact of tighter budgets and rolling windows.
Recency-weighted averages and low-order ARIMA models provide strong, stable baselines. Hybrid and robust variants often improve performance, though gains are not universal. The optimizer frequently favors a 3-5-2 shape, reflecting midfielders’ outsized contribution to FPL scoring. Overall, translating noisy forecasts into constraint-aware squads yields transparent, reproducible, and practically useful improvements.
Future work includes embedding transfers and chips in a rolling-horizon setting, allowing week-to-week captaincy, refining goalkeeper objectives, and enriching the feature set with external match-level data. Potential future changes in the FPL rules can also be incorporated into the optimization model.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
CRediT authorship contribution statement
Danial Ramezani: conceptualization, data curation, methodology, experiment, visualization, writing, editing; Tai Dinh: experiment, visualization, writing, editing, validation.