### v3

# A data-driven framework for team selection in Fantasy Premier League

Abstract
Fantasy football is a billion-dollar industry with millions of participants. Under a fixed budget, managers select squads to maximize future Fantasy Premier League (FPL) points. This study formulates lineup selection as data-driven optimization and develops deterministic and robust mixed-integer linear programs that choose the starting eleven, bench, and captain under budget, formation, and club-quota constraints (maximum three players per club). The objective is parameterized by a hybrid scoring metric that combines realized FPL points with predictions from a linear regression model trained on match-performance features identified using exploratory data analysis techniques. The study benchmarks alternative objectives and cost estimators, including simple and recency-weighted averages, exponential smoothing, autoregressive integrated moving average (ARIMA), and Monte Carlo simulation. Experiments on the 2023/24 Premier League season show that ARIMA with a constrained budget and a rolling window yields the most consistent out-of-sample performance; weighted averages and Monte Carlo are also competitive. Robust variants and hybrid scoring metrics improve some objectives but are not uniformly superior. The framework provides transparent decision support for fantasy roster construction and extends to FPL chips, multi-week rolling-horizon transfer planning, and week-by-week dynamic captaincy.
1 Introduction
Fantasy sports are a billion-dollar industry, with the market size expected to reach \$84.98 billion by 2032111https://www.skyquestt.com/report/fantasy-sports-market/. Fantasy football (soccer) is a leading segment: the official Fantasy Premier League (FPL) alone engaged over 11 million users in 2025222https://fantasy.premierleague.com/. In FPL, managers assemble a 15-player squad (two goalkeepers, five defenders, five midfielders, and three forwards) within a fixed budget of £100.0 m. Player prices evolve with performance and demand, and only the starting XI score in a given gameweek, with the captain’s points doubled; optional season-limited “chips” can alter scoring in specific weeks. Although the rules are simple, the weekly decision problem, which selects a cost-feasible starting XI and a captain under formation and club-quota constraints, with uncertain future performance, is nontrivial. With a global user base exceeding tens of millions and an industry valuation in the billions of dollars, the economic and social impact of fantasy sports cannot be overlooked. For participants, optimal decision-making translates into both competitive success and, in many cases, tangible financial rewards.
Against this backdrop, fantasy football has attracted attention across research and practitioner communities. Qualitative studies examine impacts on related markets and behavior, including merchandise sales (drayer2010effects), game attendance (nesbit2010impact), and participant well-being (wilkins2021exploring). Quantitative research has explored performance prediction and strategy design using econometric and machine learning techniques. The increasing availability of public data and analytics tooling has further lowered barriers to empirical work and large-scale evaluation in this domain.
Despite this progress, most quantitative FPL studies emphasize forecasting weekly player points, with comparatively less attention to integrating those forecasts into a decision model that captures the full set of FPL constraints and jointly selects the starting XI and captain. Decision-analytic approaches exist (matthews2012competing), but captaincy is typically handled outside the optimization pipeline (e.g., via crowds or heuristics) rather than optimized jointly (bhatt2019should). Moreover, robustness to estimation error is rarely addressed, and there is limited comparative evidence on how alternative objective functions, selection budget, rolling-window specifications, or cost-vector estimation methods affect out-of-sample outcomes within a unified optimization framework.
Accordingly, we concentrate on two practical concerns: (i) converting noisy performance forecasts into an actionable weekly decision that jointly determines the starting XI and captain under budget, formation, and club-quota constraints; and (ii) mitigating sensitivity to estimation error and week-to-week variance. We address these by coupling integer programming models (with an explicit captaincy variable and operational constraints) with data-driven cost estimation and a hybrid scoring metric, and by evaluating a robust variant designed to hedge against misspecification. We also studied how budget constraints can change the results.
Generally, this paper makes the following contributions:
  • We formulate FPL lineup selection as a data-driven optimization problem and develop deterministic and robust integer-programming models that choose the starting XI, bench players, and captain under budget, formation, and club-quota constraints.
  • We propose a hybrid scoring metric that combines realized FPL points with predictions from a ridge regression model trained on interpretable match-performance features.
  • We build a unified pipeline that benchmarks alternative objectives and multiple forecasting approaches (including recency-weighted averaging, low-order ARIMA models, and simulation-based estimators) and provide an empirical assessment on the 2023/24 Premier League season. This analysis includes a rolling-window specification and examines how constraining the budget for the starting XI affects out-of-sample performance, yielding managerial insights into emergent formations and player selection patterns. To the best of our knowledge, this is among the first systematic studies of data-driven team selection in FPL that leverages a hybrid predictive metric within a robust decision-making framework and provides such comprehensive analysis.
  • We publicly release all datasets, preprocessing scripts, and experiment notebooks in a GitHub repository [^3] https://github.com/DanialRamezani/Data-Driven-FPL to facilitate reproducibility and encourage further research and practical adoption.
The remainder of the paper is organized as follows. Sec. 2 reviews related work. Sec. 3 presents preliminaries and the optimization models. Sec. 4 introduces the proposed framework. Sec. 5 details the cost-vector estimation strategies and alternative objectives. Sec. 6 describes the dataset and feature construction. Sec. 7 shows exploratory data analysis results. Sec. 8 reports the empirical results and discusses the optimal lineups and formations. Sec. 9 concludes and outlines directions for future research.
2 Related work
Several strands of prior work relate to our study (ati2024using). First, decision-analytic and optimization approaches frame fantasy team management as an explicit choice problem. matthews2012competing model the FPL task as a belief-state Markov decision process with a tractable Bayesian Q-learning algorithm, casting weekly choices as a multi-dimensional knapsack and achieving around top-percentile performance against millions of human players. Extending optimization ideas to the NFL, becker2016analytical formulate a mixed-integer program to support draft selections and weekly lineup management using historical data and predictions. In FPL, gupta2019time propose a linear selection model for squad construction; while intentionally simple and omitting some operational constraints (e.g., optimal formation), they pair it with a hybrid RNN–ARIMA forecaster to supply objective inputs. More recently, metaheuristics and integrated pipelines have appeared: aribo2024machine apply a genetic algorithm for team selection and budget allocation, and venter2024optimisation combine optimization with forecasting in a practical system that attained a top-$4\%$ rank in the 2021/22 season. Taken together, these studies underscore the promise of explicit decision models, though they differ in how fully they encode FPL rules and whether captaincy is optimized jointly with the starting XI.
A second line of research develops predictive models for player performance, often treating the decision task separately. For NFL quarterbacks, lutz2015fantasy evaluated neural networks and support vector machines to improve score prediction, and steenkiste2015finding compared linear regression, random forests, and multivariate adaptive regression splines for next-week outcomes. Hybrid time-series ideas enter the literature with gupta2019time, who fuse a recurrent neural network with ARIMA to estimate the cost vector used downstream. Within FPL, bangdiwala2022using assessed several machine-learning models and reported linear regression as the best performer on their dataset, while pokharel2022fantasy used an XGBoost regressor to study selection and transfer strategies through a return-on-investment lens and examine the effects of midweek cup fixtures. Deep sequence models have also been explored: lombu2024predicting compared convolutional neural networks with long short-term memory networks using form from the previous five matches, finding LSTM superior. Overall, the forecasting literature is methodologically rich, yet downstream optimization is often decoupled from prediction.
Complementing model-based prediction, researchers have examined crowd signals and ancillary tasks that influence managerial choices and market dynamics. bhatt2019should analyzed Twitter data for captain selection and found that the “wisdom of the crowd” outperforms expert analyst recommendations in their setting. In a related operational facet, khamsan2019handling addressed highly imbalanced classification when predicting virtual player price changes in FPL, where the majority class corresponds to no change.
Positioning this study within the literature, we bridge these strands by integrating data-driven forecasting with an integer optimization model that jointly selects the starting XI and the captain under full FPL constraints, and by assessing robustness to estimation error within the same framework. This unified treatment enables a comparative evaluation of objective functions and cost-vector estimators, linking predictive accuracy to operational decision quality.
3 Preliminaries and mathematical formulations
An FPL squad consists of 15 players: two goalkeepers (GK), five defenders (DEF), five midfielders (MID), and three forwards (FWD). Each gameweek, only the starting XI scores; one of the XI is named captain and receives double points each week. At most three players may come from any single Premier League team, and the total squad value cannot exceed £100.
For simplicity, gamers who select football players will be referred to as team owners, while professional Premier League footballers are called players. As previously mentioned, the goal of this paper is to select players in a way that maximizes the total points in the upcoming weeks. Therefore, special and one-time cards are not taken into account. The selection of the starting eleven can be formulated as the following integer programming model:
  • $n$: total number of players.
  • $c_{j}$: expected points for player $j$.
  • $b$: total budget for the starting eleven.
  • $v_{j}$: value (price) of player $j$.
  • $\text{min\_limit}_{k}$: minimum number of players to be selected from position $k$.
  • $\text{max\_limit}_{k}$: maximum number of players to be selected from position $k$.
  • $\text{Team}_{t}$: set of players from team $t$.
Decision variables:
  • $x_{j}\in\{0,1\}$ for $j=1,2,\dots,n$: $x_{j}=1$ if player $j$ is selected in the starting eleven; 0 otherwise.
  • $y_{j}\in\{0,1\}$ for $j=1,2,\dots,n$: $y_{j}=1$ if player $j$ is selected as captain; 0 otherwise.
Objective:
$\max\;Z=\sum_{j=1}^{n}c_{j}x_{j}+\sum_{j=1}^{n}c_{j}y_{j}$ (1)
The objective maximizes the expected reward of the team. In FPL, the captain’s score is doubled; modeling captaincy with $y_{j}$ adds a second $c_{j}$ term when player $j$ is captain.
Subject to:
$\sum_{j=1}^{n}x_{j}=11$ (2)
Eq. 2 indicates that exactly eleven starters are selected.
$\sum_{j=1}^{n}v_{j}x_{j}\leq b$ (3)
Eq. 3 indicates that the total cost of the starting eleven does not exceed the budget $b$.
$\sum_{j=1}^{n}y_{j}=1$ (4)
Eq. 4 indicates that exactly one captain is chosen.
$y_{j}\leq x_{j},\quad\forall j=1,2,\dots,n$ (5)
Eq. 5 indicates that the captain must be among the selected starters.
$\sum_{j\in\text{Position}_{k}}x_{j}\geq\text{min\_limit}_{k},\quad\forall k\in\{\text{GK},\text{DEF},\text{MID},\text{FWD}\}$ (6)
Eq. 6 enforces the minimum positional counts for a legal formation.
$\sum_{j\in\text{Position}_{k}}x_{j}\leq\text{max\_limit}_{k},\quad\forall k\in\{\text{GK},\text{DEF},\text{MID},\text{FWD}\}$ (7)
Eq. 7 enforces the maximum positional counts for a legal formation.
$\sum_{j\in\text{Team}_{t}}x_{j}\leq 3,\quad\forall t=1,2,\dots,20$ (8)
Eq. 8 ensures that at most three players may be selected from any single Premier League club.
In this work, various methods are analyzed to estimate the best choice of $c_{j}$ for out-of-sample performance. Although decision-makers are free to define any budget below 100, in this study, we set $b=83.5$ as a default value. The rationale is straightforward: the model must also account for four reserve players, and in FPL the cheapest players typically cost around four million pounds each, totaling 16 million. An additional 0.5 million is added as a buffer to prevent budget violations. Thus, the formulation focuses on selecting the best fixed starting squad, while team owners are free to choose four inexpensive reserve players for a total of up to £ 16.5 million.
3.1 Optimizing bench
In the official FPL, managers register 15 players with positional totals $2\text{--}5\text{--}5\text{--}3$ (two goalkeepers, five defenders, five midfielders, three forwards). We therefore optimize the four bench slots subject to budget, team quota, and positional complement constraints so that the combined roster (starters + bench) satisfies the official totals.
Additional decision variables:
  • $x^{b}_{j}\in\{0,1\}$ for $j=1,2,\dots,n$: $x^{b}_{j}=1$ if player $j$ is selected for the bench; 0 otherwise.
Additional parameters:
  • $b^{\text{bench}}$: budget available for the bench (100b100-b).
  • $r_{k}$: required number of players in position $k$ for the bench.
  • $s_{t}$: number of players from team $t$ already selected in the starting XI.
  • $S$: set of players already selected in the starting XI.
Bench objective:
$\max\;Z^{bench}=\sum_{j=1}^{n}c_{j}x^{b}_{j}$ (9)
Eq. 9 maximizes the expected contribution of bench players, providing insurance when a starter does not feature.
Bench constraints:
$\sum_{j\in\text{Position}_{k}}x^{b}_{j}=r_{k},\quad\forall k\in\{\text{GK},\text{DEF},\text{MID},\text{FWD}\}$ (10)
Eq. 10 enforces positional requirements for the bench so the full 15-man roster satisfies FPL’s 2–5–5–3 totals.
$\sum_{j=1}^{n}v_{j}x^{b}_{j}\leq b^{\text{bench}}$ (11)
Eq. 11 keeps the bench within the remaining budget $b^{\text{bench}}=100-b$.
$\sum_{j\in\text{Team}_{t}}x^{b}_{j}+s_{t}\leq 3,\quad\forall t=1,2,\dots,20$ (12)
Eq. 12 extends the three-player-per-team restriction to the full squad (starters + bench).
$x^{b}_{j}=0,\quad\forall j\in S$ (13)
Eq. 13 prevents duplication: no player may appear in both the starting XI and the bench.
3.2 Robust optimization
Robust optimization is a decision-making framework under uncertainty. Unlike stochastic programming, which requires perfect knowledge of the distribution of the uncertain parameters and scenario generation, robust optimization focuses on ensuring that the model solution remains feasible and performs well under the worst-case realization of uncertainty.
Let the actual expected points be uncertain and lie in a box uncertainty set:
$\mathcal{U}=\left\{c\in\mathbb{R}^{n}:c_{j}\in[\bar{c}_{j}-d_{j},\bar{c}_{j}+d_{j}]\;\;\forall j\right\},$ (14)
In Eq. 14, $\bar{c}_{j}$ is the nominal expected score and $d_{j}\geq 0$ is the uncertainty margin for player $j$.
Finally, the worst-case objective is formulated as follows:
$\max_{x,y}\;\min_{c\in\mathcal{U}}\sum_{j=1}^{n}c_{j}(x_{j}+y_{j}).$ (15)
Eq. 15 ensures that the selected team maximizes the worst-case total score, making the solution robust against possible overestimation in the expected scores and week-to-week variance.
Figure 1: The proposed framework. Modeling produces expected points, which feed the optimization to select XI and captain under FPL constraints.
4 The proposed framework
Figure 1 shows the proposed end-to-end pipeline from the merged FPL dataset to a team recommendation. We begin with a player–gameweek (GW) panel, verify types, remove duplicates, and guard against leakage by excluding any same-GW outcomes (e.g., bonus/BPS) from predictors. We fix a train→test split (GW1–26 → GW27–38) and treat missing player–GW rows as unavailability rather than an explicit injury flag. For the target GW we also built a compact table schema for the optimizer—name, position, team, value, and a slot for expected points—with the slot filled by the models later; modeling inputs are standardized.
[…]
The modeling stage fits five families: (1) averages/smoothing, (2) ARIMA variants, (3) linear trend regression, (4) simulation (bootstrap and parametric), and (5) the hybrid that combines recent outcomes with the Ridge feature model. These produce per-player expected points $c_{j}$ (and, when applicable, uncertainty widths $d_{j}$).
Finally, a 0–1 integer program selects the starting XI and captain under FPL rules-budget, valid formations, and at most three players per club-with the captain required to start. A bench is then chosen from the remaining budget to complete a 2–5–5–3 squad under the same club quota. Besides maximizing expected points, we also report surrogate XI objectives (xGI-xGC or ICT) and an optional robust variant that penalizes scores by uncertainty ($c_{j}\pm d_{j}$). We output the recommended XI, captain, and bench, along with formation trends, weekly and cumulative points, method leaderboards, and budget/rolling-window sensitivity analyses.
5 Estimating cost vectors
This section details how we estimate the per–player expected points $c_{j}$ used by the optimization models in Sec. 3. All methods consume historical weekly points up to a fixed split week and return a single scalar summary per player. Let the season length be $N=38$ gameweeks and the train/test split occur at gameweek $\tau$ (train: weeks $1{:}\tau$; test: weeks $\tau{+}1{:}N$). For a player with time series $\{p_{1},\dots,p_{\tau}\}$, the goal is to produce a representative value $c$ summarizing expected future performance.
5.1 Averaging-based estimators
We begin with simple summaries of $\{p_{t}\}$.
Simple average
$\widehat{c}^{\,\text{SA}}\;=\;\frac{1}{\tau}\sum_{t=1}^{\tau}p_{t}$ (16)
Eq. 16 treats all observed weeks equally.
Weighted average (recency emphasis)
Let the linear weights be $w_{t}=t\big/\sum_{i=1}^{\tau}i$.
$\widehat{c}^{\,\text{WA}}\;=\;\sum_{t=1}^{\tau}w_{t}\,p_{t}$ (17)
Eq. 17 indicates that more recent weeks receive larger weights; week $\tau$ has the highest weight.
Exponential smoothing (Holt’s linear trend)
We fit a level–trend model (no seasonality):
$\hat{p}_{t}$ $=\ell_{t-1}+b_{t-1}$ (18)
$\ell_{t}$ $=\alpha\,p_{t}+(1-\alpha)(\ell_{t-1}+b_{t-1})$ (19)
$b_{t}$ $=\beta\,(\ell_{t}-\ell_{t-1})+(1-\beta)b_{t-1}$ (20)
In the above equations, $\ell_{t}$ and $b_{t}$ denote level and trend; $\alpha,\beta\in[0,1]$ are estimated from data. Forecast the remaining weeks $\{\hat{p}_{\tau+1},\dots,\hat{p}_{N}\}$ and summarize by
$\widehat{c}^{\,\text{ES}}\;=\;\frac{1}{N-\tau}\sum_{k=1}^{N-\tau}\hat{p}_{\tau+k}.$ (21)
Eq. 21 averages the future forecasts to produce a single score.
5.2 Simulation-based estimators
We approximate future outcomes by sampling from historical behavior.
Bootstrap (nonparametric)
Draw $B$ resamples of length $N-\tau$ with replacement from $\{p_{1},\dots,p_{\tau}\}$; average within resamples, then across resamples:
$\widehat{c}^{\,\text{BOOT}}\;=\;\frac{1}{B}\sum_{b=1}^{B}\left(\frac{1}{N-\tau}\sum_{k=1}^{N-\tau}p^{*(b)}_{k}\right)$ (22)
Eq. 22 makes no distributional assumptions; $p^{*(b)}_{k}$ are bootstrap draws.
Monte Carlo (parametric)
Assume $p_{t}$ are i.i.d. with mean $\bar{p}=\frac{1}{\tau}\sum_{t=1}^{\tau}p_{t}$ and variance $s^{2}$. Simulate $B$ future paths from a chosen family (e.g., Gaussian or truncated Gaussian) with parameters $(\bar{p},s^{2})$ and compute the same summary as in (22) to obtain $\widehat{c}^{\,\text{MC}}$.
5.3 Autoregressive integrated moving average (ARIMA)
We fit an $\text{ARIMA}(p,d,q)$ model to $\{p_{t}\}_{t=1}^{\tau}$:
$\phi(B)\,(1-B)^{d}\,p_{t}\;=\;\theta(B)\,\varepsilon_{t}$ (23)
In Eq. 23, $B$ is the backshift operator; $\phi(\cdot)$ and $\theta(\cdot)$ are AR and MA polynomials; $\varepsilon_{t}$ is white noise. Rolling forecasts $\{\hat{p}_{\tau+1}$, $\dots$, $\hat{p}_{N}\}$ are summarized by
$\widehat{c}^{\,\text{ARIMA}}\;=\;\frac{1}{N-\tau}\sum_{k=1}^{N-\tau}\hat{p}_{\tau+k}.$ (24)
Eq. 24 uses the average of multi-step forecasts as a single cost value.
5.4 Linear regression (time-on-trend)
Regress points on week index $w_{t}=t$ for $t=1,\dots,\tau$:
$p_{t}\;=\;\beta_{0}+\beta_{1}w_{t}+\varepsilon_{t}$ (25)
In Eq. 25, $\beta_{0},\beta_{1}$ are OLS estimates; $\varepsilon_{t}$ are residuals. Predict weeks $w_{\tau+1},\dots,w_{N}$ and summarize:
$\widehat{c}^{\,\text{LR}}\;=\;\frac{1}{N-\tau}\sum_{k=\tau+1}^{N}\hat{p}_{k}.$ (26)
Eq. 26 averages the linear-trend forecasts to a single score.
5.5 Hybrid method (ridge + realized points)
The hybrid approach augments weekly points with match-performance features that are not directly scored by FPL but correlate with future returns (e.g., ICT components, starts, shots). Let $\mathbf{X}$ be feature vectors and $\mathbf{y}$ be targets (weekly points) up to week $\tau$. Standardize features and fit ridge regression:
$\hat{\mathbf{w}}\;=\;\arg\min_{\mathbf{w}}\left\{\left\|\mathbf{y}-\mathbf{Xw}\right\|_{2}^{2}+\alpha\left\|\mathbf{w}\right\|_{2}^{2}\right\}$ (27)
In Eq. 27, $\alpha\geq 0$ controls shrinkage. Use the fitted model to predict per–player season-forward means $\widehat{y}$ (e.g., by averaging predicted weeks $\tau{+}1{:}N$). Combine normalized realized and predicted values:
$\widehat{c}^{\,\text{HYB}}\;=\;(1-\lambda)\,y^{\text{norm}}\;+\;\lambda\,\widehat{y}^{\,\text{norm}},\quad\lambda\in[0,1].$ (28)
In Eq. 28, $y^{\text{norm}}$ and $\widehat{y}^{\text{norm}}$ are per–player normalizations (e.g., z scores or min–max). We evaluate $\lambda=\tfrac{1}{3}$ (2:1 favoring realized) and $\lambda=\tfrac{2}{3}$ (1:2 favoring predicted).
5.6 Alternative objective surrogates
For completeness, we also replace points with proxy targets.
ICT index maximization
$\max\sum_{j=1}^{n}\text{ICT}_{j}\,(x_{j}+y_{j})$ (29)
In Eq. 29, $\text{ICT}_{j}$ denotes the expected ICT score of player $j$; the captain contribution is doubled via $y_{j}$.
Attack–defense trade-off (EGI-EGC)
$\max\sum_{j=1}^{n}\big(\text{EGI}_{j}-\text{EGC}_{j}\big)\,(x_{j}+y_{j})$ (30)
In Eq. 30, $\text{EGI}_{j}$ is expected goal involvement; $\text{EGC}_{j}$ is expected goals conceded for player $j$. Higher values prefer attacking contributors while discouraging players expected to concede.
6 Dataset and experimental setup
We use the publicly available FPL dataset [^4] https://github.com/vaastav/Fantasy-Premier-League/, which provides a player–gameweek panel for each Premier League season (38 gameweeks). FPL awards points according to role-specific rules (e.g., clean sheets for defenders/goalkeepers; goals/assists for attackers), while some events apply to all positions (e.g., minutes played, cards) [^5] https://www.premierleague.com/news/2174909/. Feature names follow the official FPL statistics glossary [^6] https://fantasy.premierleague.com/statistics/.
The unit of analysis is a player–gameweek. Unless stated otherwise, we use the 2023/24 season with $N=38$ gameweeks and a fixed split at gameweek $\tau$: training on weeks $1{:}\tau$ and evaluating on weeks $\tau{+}1{:}N$ (consistent with Sec. 5). Player identity is the FPL unique player code; club membership is the contemporaneous club in that gameweek (relevant for the club–quota constraints in Sec. 3). Prices are scaled to millions, as in our code (value divided by 10), and are used downstream for the budget constraint.
Our outcome is gameweek total points, denoted $p_{t}$ in Sec. 5. To avoid target leakage, we exclude contemporaneous outcomes and referee events as predictors for all models: goals scored, clean sheets, goals conceded, penalties missed, saves, penalties saved, own goals, yellow cards, red cards, assists, team scores (team h score, team a score), and the bonus system (bps, bonus). We also drop schedule/row identifiers that do not encode player features (opponent team, GW/round, kickoff time, fixture, was home) as modeling features. The target–week minutes is never used as a predictor. The resulting filtered table is used to construct the covariates for the hybrid/surrogate experiments below, while prices (values), teams, and positions are retained for the optimizer.
The following shows the detailed setup of the prediction models.
  1. (a)
    History-only forecasters include simple and weighted averages, exponential smoothing, ARIMA, linear trend, and simulation (bootstrap/normal Monte Carlo). These methods estimate each player’s expected points using only $\{p_{1},\ldots,p_{\tau}\}$; no exogenous features are consumed. We also retain the empirical standard deviation of past $p_{t}$ as a simple uncertainty proxy.
  2. (b)
    We form per–player aggregates over the training window for non–scoring, pre–deadline proxies that survive filtering: ict index,expected goals (xG), expected assists (xA), expected goal involvements (xGI), expected goals conceded (xGC), selected, starts.
    • For each position, we fit a ridge regression of
      total points on the proxies above (standardized), and combine the ridge prediction with the normalized historical mean of total points to obtain a hybrid score.
    • For interpretability, we also fit an XGBoost regressor on the same proxies and use SHAP library to visualize influential features. This step is diagnostic only and is not fed to the optimizer in the reported runs.
    • Surrogates: maximize $\mathrm{EGI}-\mathrm{xGC}$ and use a robust ICT score $\emph{ict index}-\mathrm{sd}(\emph{ict index})$.
Unless stated otherwise, the integer program in Sec. 3 uses value for the budget, position for formation limits, and team for the $\leq 3$ per–club rule. The starting XI objective is chosen from the set above (expected points, hybrid score, EGI-xGC, robust ICT). The bench is ordered by the history-only expected points.
7 Exploratory data analysis
Figure 2: Correlation between selected features and total points by positions.
Figure 3: Pearson correlation heatmap of modeling features.
Figure 4: RainCloud plots of top numeric features by position.
(a) Goalkeepers
(b) Defenders
(c) Midfielders
(d) Forwards
Figure 5: SHAP value summaries for key features by position.
Figure 6: Hierarchical clustering of strategy behaviors. The heatmap reports the absolute Spearman correlation between weekly team scores for all strategies, and the dendrogram groups runs with similar week-to-week profiles.
Fig. 2 shows Pearson correlations between the modeling features and gameweek total points by position, computed on the training window ($\mathrm{GW}\leq\tau$). Across roles, ict index is the strongest single proxy, and exposure signals such as starts (and selected) also correlate positively with points. For attackers (Forward, Midfielder), expected goal involvements and expected goals exhibit sizeable positive correlations; these signals are weaker but still positive for Defenders. For the Goal Keeper panel, starts and ict index dominate, with other proxies contributing modestly.
Fig. 3 highlights a clear “attacking” cluster. Features including xG, xA, xGI, and ict index are all strongly and mutually correlated, and each shows a solid positive relationship with total points. Starts and selected also track positively with points (proxies for minutes/availability), while value co-moves with performance and popularity. By contrast, xGC has the weakest—and likely slightly negative—association with total points and only mild links to the attacking features. Overall, the tight correlations among xG/xA/xGI/ICT indicate notable multicollinearity, so using regularization or a reduced set (e.g., keep xGI) would avoid redundancy.
Fig. 4 shows zero-inflation and heavy right tails for attacking metrics (xA, xG, xGI, ICT), position-dependent heteroscedasticity (GK/DEF tight; MID/FWD wide), and strong collinearity among attacking features, with value rising along the same gradient. This shape favors recency-aware, shrinkage forecasts with the rolling ARIMA and weighted averages because they 1) stabilize means in the presence of outliers/hauls, 2) adapt to non-stationary form and changing minutes, and 3) avoid exploding variance that simple means or Monte-Carlo draws suffer in heavy-tailed data. Collinearity among xG/xA/xGI/ICT motivates feature parsimony or ridge in the hybrid metric (often collapsing to xGI + regularization). On the optimization side, the distributions justify a budget- and formation-constrained IP: it allocates spend and captaincy toward the higher-mean, higher-upside MID/FWD slots while using cheaper, lower-variance GK/DEF to meet minima; an objective that rewards xGI and/or penalizes xGC aligns with the positional patterns seen.
Fig. 5 shows position-wise SHAP summaries from the diagnostic XGBoost: the horizontal spread indicates importance (wider = larger impact), and color encodes the raw feature value (pink = high, blue = low). It can be observed that the ICT index is the most influential predictor across DEF/MID/FWD, with higher ICT consistently pushing predictions up. For attackers, xGI (and its components xG, xA) strongly and positively drive the model, while starts mainly captures availability effects. On the defensive side, expected goals conceded (xGC) has a clear negative contribution (higher xGC leads to a lower output), especially for defenders, and starts is most important for goalkeepers. Selected% and value add comparatively little marginal signal once the other features are included.
8 Experimental results and discussion
Figure 6 summarizes how all strategies move week-to-week. Each cell is the absolute Spearman correlation between two methods’ weekly team scores (GW27-38), and the dendrogram clusters methods with similar temporal profiles. Clear block-diagonal structure appears: rolling ARIMA variants form a tight cluster, simple/weighted/ExpSm averages cluster together, simulation/Monte-Carlo runs form another block, and ICT/linear-regression-based strategies sit farther away with weaker affinities to the averaging/ARIMA families. Budgets of the same method line up next to one another (e.g., £80-£65), indicating that the budget mainly shifts level but not the week-to-week pattern. Hybrid methods typically lie between their parent families (e.g., Hybrid Weighted near Weighted Avg.), while robust versions are slightly less correlated, reflecting their penalty on uncertainty. Overall, the map shows substantial redundancy—many methods behave very similarly—so in the remainder we focus on a representative subset and benchmark uplifts relative to a single strong baseline.
(a) Averaging methods and robust counterparts
(b) Simulation methods (bootstrap vs. Monte Carlo)
Figure 7: Out-of-sample performance: averaging and simulation families.
Figure 8: Out-of-sample performance of ARIMA variants. Simpler models (ARIMA (0,0,1) and ARIMA (1,0,1)) outperform higher-order configurations.
Figure 9: Alternative objectives: ICT, robust ICT, and involvement (EGI-EGC).
(a) Hybrid over averaging methods
(b) Hybrid over simulation methods
Figure 10: Effect of the hybrid metric on averaging and simulation families.
(a) Hybrid vs. ARIMA (0,0,1)
(b) Hybrid ICT vs. ICT and linear regression
Figure 11: Hybridization applied to ARIMA and ICT/linear regression.
Having analyzed the similarity of strategy behaviors (Figure 3), we now focus on the models that achieve the strongest out-of-sample performance. We evaluate strategies by cumulative team points, training on gameweeks 1-26 of the 2023/24 season and evaluating on gameweeks 27-38. In gameweek 29 many player records are missing; instead of discarding teams, we apply the bench-substitution policy described earlier. All subsequent plots report cumulative team points (with no chips played). To avoid clutter and highlight practical takeaways, we show only the best-performing variant within each method family in the remaining figures and tables.
Figure 7 reports cumulative out-of-sample points for the averaging and simulation families. Panel (a) shows that the Weighted Avg. (recency-weighted mean) clearly dominates Simple Avg. and Exponential Smoothing over GW 27–38, suggesting that putting more weight on recent form is beneficial for FPL scoring. The robust versions of these methods track their non-robust counterparts from below: penalizing uncertainty mainly trims upside weeks without delivering visible protection in bad weeks, so robust curves never catch up. Panel (b) compares Simulation (Non Parametric) (bootstrap) and Monte Carlo Simulation to the simple average baseline; both simulation curves almost coincide with each other and with Simple Avg., indicating that, at this horizon and sample size, adding a stochastic layer (either empirical or parametric) offers little advantage over a deterministic average of past points.
Figure 8 compares nine ARIMA specifications. The simplest models without differencing, ARIMA(0,0,1) and ARIMA(1,0,1), clearly dominate in this static setting, finishing the horizon at roughly 680 and 660 cumulative points, respectively. ARIMA(1,0,0) forms a second tier, ending about 50 points behind, while all higher-order or differenced variants (e.g., ARIMA(1,1,1), ARIMA(2,1,1), ARIMA(2,1,2)) lag steadily throughout GW 27–38. This pattern is consistent with the short, noisy weekly horizon: lightly parameterized ARIMA models capture the main signal, whereas more complex configurations tend to overfit in-sample fluctuations that do not persist out of sample. However, when ARIMA processes are embedded in our rolling-selection framework with a constrained starting-XI budget, ARIMA(1,0,0) proves more stable and ultimately delivers the best overall performance; for this reason, ARIMA(1,0,0) with a £70m XI budget is adopted as the benchmark ARIMA model in our later cross-family and external comparisons.
Figure 9 compares alternative objectives: ICT, robust ICT, and an attack-defense proxy based on maximizing expected goal involvement minus expected goals conceded (EGI-EGC). It is evident that robust ICT marginally outperforms deterministic ICT, continuing the trend seen in the averaging family, though the improvement is modest. On the other hand, the involvement proxy underperforms significantly, likely due to the inability to capture indirect or additional important features such as those related to team quality, minutes played, and so on.
We next study the effect of the hybrid scoring metric in Eq. (28). Figure 10 applies it to the averaging and simulation families, while Figure 11 applies it to ARIMA, ICT, and linear regression. Across all families, the 1:2 setting (placing two-thirds of the weight on the model-based score and one-third on realised FPL points) is the most reliable: it typically sits above or on top of the original method’s curve, delivering a modest but consistent uplift. In contrast, the 2:1 setting often underperforms the base method, suggesting that overweighting raw historical points re-introduces the very noise the predictive model is designed to filter out.
For the stronger baselines—ARIMA (0,0,1), ICT, and linear regression—the same pattern holds: the hybrid 1:2 variants track or exceed the non-hybrid trajectories over most of the horizon, whereas the 2:1 versions are uniformly worse. This indicates that hybridization is most effective when it uses realized points as a light calibration signal on top of a stable predictive score, rather than as the dominant driver of team selection.
Figure 12: Budget sensitivity (static teams): Simple average.
The base budget is our default b = £83.5 for the starting XI (with the total FPL budget still at £100). We then vary this cap between £55 and £80 to study sensitivity. Figure 12 shows how different budget caps affect a single static team built with the simple‐average predictor. The bump chart tracks the rank (1 = best) of each budget’s cumulative score from GW 27 to 38, with final totals reported on the right. After some early crossings, the ranking stabilises: the £80 budget quickly takes over and remains first to the end, followed by £75, while the nominal “Base” budget finishes only fourth behind £70. The two lowest caps (£60 and especially £55) stay near the bottom throughout. Overall, the figure suggests clear but non-linear budget sensitivity: moving from very low to moderate budgets yields large gains, but simply increasing the cap beyond about £75 does not guarantee better static performance.
Figure 13: Budget sensitivity (static teams): Weighted average.
Figure 13 illustrates the budget sensitivity for static teams using the weighted average method. The plot shows how teams with different budgets (ranging from £80 to £60) perform across the gameweeks. It is evident that teams with a higher budget consistently rank better, with the £80 budget achieving the highest rank and maintaining strong performance throughout the season. Conversely, teams with lower budgets, particularly £60, underperform, indicating that a more substantial budget enables better team selection and performance. This reinforces the importance of budget flexibility in achieving competitive rankings, as the weighted average method benefits from selecting players with higher recent performance, which requires greater financial resources.
Figure 14: Rolling selection: ARIMA (1,0,0).
Figure 14 illustrates the budget sensitivity and performance of static and rolling selection strategies using ARIMA (1,0,0). The plot compares static teams with varying budgets (£80, £75, £70, £65, £60) and rolling teams for each corresponding budget. While static teams with higher budgets generally perform better, the rolling selection strategy (denoted by the "Roll" lines) exhibits competitive performance, especially for the £75 and £70 budgets, which outperform the static £80 variant during some gameweeks. This highlights that rolling selection can adapt dynamically to ongoing changes and might offer an edge over static selection in certain budget ranges, even if it has higher variance in rankings.
Figure 15: Rolling selection: ARIMA (1,0,1).
Figure 15 shows the budget–rank trajectories for static and rolling teams under ARIMA (1,0,1). Rolling selection is again highly competitive: the £80 and £70 rolling variants dominate the ranking from roughly GW 31 onward and finish the season as the top two configurations. The static base and £80–£75 teams form a second tier, while low-budget static teams (£65–£60) remain near the bottom throughout. The £60 rolling team still outperforms its static £60 counterpart, indicating that dynamic re-optimization helps even when resources are tight, although the benefit is most pronounced at higher budgets.
Figure 16: Budget sensitivity (static teams): ARIMA (0,0,1).
Figure 16 summarizes budget sensitivity for static ARIMA (0,0,1) teams. The coloured strip at the top shows which budget produced the highest weekly score in each gameweek (GW 27–38), while the lollipops on the bottom panel report the GW 38 scores for each budget together with the number of weekly “wins” (W) in the legend. The £75 variant records the most weekly wins (4), followed by the base-budget squad (3) and the £80 and £65 variants (2 each), indicating that mid-range budgets frequently top the table in individual weeks. However, the base team still attains the highest final-week score (41 points), with £75 close behind and the low-budget £65 side clearly lagging. Overall, ARIMA (0,0,1) exhibits only modest sensitivity to budget: extra funds offer some benefit at the extremes, but the performance gap between the base and mid-range budgets is small relative to week-to-week volatility.
Figure 17: Budget sensitivity (static teams): ARIMA (1,0,0).
Figure 17 reports budget sensitivity for static ARIMA (1,0,0) teams. The top strip shows which budget wins each gameweek; most tiles are orange or green, confirming that the £80 and £75 variants dominate weekly leadership (four wins each), while the base and cheaper budgets only top the table once. The lollipops in the lower panel summarize GW 38 performance: the £80 squad achieves the highest final score (49 points), followed by £75 (43) and the base budget (39), with £70, £60, and £65 trailing more clearly. Overall, ARIMA (1,0,0) benefits more from extra budget than ARIMA (0,0,1): mid-to-high budgets (£75–£80) deliver both more weekly wins and a noticeable improvement in final-week performance over the base case.
Figure 18: Budget sensitivity (static teams): ARIMA (1,0,1).
Figure 18 shows budget sensitivity for static ARIMA (1,0,1) teams. The top winner strip indicates which budget leads in each gameweek: the £75 configuration is the most frequent weekly leader (five wins), while the base, £70, and £60 budgets each top the table in only two weeks, and £80 wins once; £65 never leads. The lollipop plot summarizes performance in the final gameweek: the £80 and base squads achieve the highest totals (55 and 54 points), followed by £75 (48 points), whereas the cheaper budgets (£70, £65, £60) cluster around the mid-20s. Overall, ARIMA (1,0,1) still rewards higher budgets, but the £75 squad offers a good compromise, capturing most weekly wins while using less budget than the full £80 baseline.
Figure 19: Budget sensitivity (static teams): ICT objective.
Figure 20: Budget sensitivity (static teams): Hybrid ICT (1:2).
Figure 19 reports budget sensitivity for the ICT-based objective with static squads. The winner strip along the top shows which budget dominates each gameweek: the base configuration wins most often (5 weeks), followed by the £80 and £70 squads (3 and 2 wins, respectively), while the £75 team wins in two late gameweeks. The lollipop panel summarizes season-end performance. Here, the £75 team achieves the highest final weekly total ($\approx$ 59 points), with the base and £80 squads essentially tied just below 50 points, and the £70 budget clearly lagging (46 points). Thus, for ICT the mid-range £75 budget offers the best overall outcome, while spending up to £80 yields little extra value and cutting to £70 is noticeably harmful.
Figure 20 shows budget sensitivity for the Hybrid ICT (1:2) objective with static squads. The winner strip indicates that leadership is more evenly shared than for pure ICT: the £80 and £70 squads each top several gameweeks (three apiece), while the base, £75, and £65 teams each lead twice. The lollipop panel summarizes the final weekly totals. Here the base-budget hybrid team again finishes first ($\approx$43 points), followed by £80 and £70 ($\approx$38 and 36 points), with £65 and £75 clearly trailing in the mid-20s. Thus, under Hybrid ICT, modest budget increases above the baseline offer only limited gains, whereas pushing the budget down to £65–£75 substantially weakens the squad, suggesting that the hybrid objective extracts most of its value from model structure rather than additional spending.
Figure 21: Budget sensitivity (static teams): Monte Carlo.
Figure 22: Rolling selection: Simple average.
Figure 21 shows how the Monte Carlo–based static teams react to budget changes. The winner strip indicates that the full-budget baseline and the £80 variant share most of the weekly wins, with the £75 team occasionally coming out on top and the cheaper £70–£60 squads almost never leading in any gameweek. The lollipop plot confirms this pattern: the baseline and £80 teams finish with very similar final weekly totals, £75 is moderately behind, and performance drops sharply once the budget falls to £70 or below. Overall, the Monte Carlo objective benefits from reasonable spending but is relatively flat between the base and £80 budgets, while aggressive under-spending severely limits its upside.
Figure 22 summarizes how the rolling Simple Avg. strategy reacts to budget changes. The winner strip shows that the £80 variant wins the largest number of individual gameweeks (five), followed by the baseline budget (four), while the £70 side wins only three weeks and the £65 team never tops the weekly ranking. The lollipop plot at the bottom reinforces this ordering: the £80 squad achieves the highest final weekly total, the baseline team is a close second, and performance drops sharply once the budget is cut to £70–£65. Overall, the rolling simple-average strategy benefits from modest extra spending but is quite fragile under tighter budget constraints.
Figure 23: Rolling selection: Weighted average.
Figure 24: Rolling selection: ARIMA (0,0,1).
Figure 23 shows budget sensitivity for the rolling Weighted Avg. strategy. The winner strip at the top indicates that no single budget dominates week by week: the £80 and £60 variants each top four and three gameweeks respectively, while the baseline budget also wins three weeks and the £70 side only two. However, the lollipop chart reveals that the baseline squad ultimately converts these weekly results into the strongest season total (42 points in the final week), with the £80 team slightly behind and the £70 and £60 teams trailing more clearly. Thus, under rolling re-optimization, the recency-weighted strategy is fairly competitive across budgets, but deep cuts to £60 still incur a noticeable loss in final performance.
Figure 24 (rolling selection with ARIMA(0,0,1)) shows that the rolling ARIMA variant again favors less-constrained budgets: the baseline squad attains the highest final weekly total (annotated on the plot $\approx$41), followed by the £70 team ($\approx$34). The £80 and £60 budgets finish noticeably lower ($\approx$26 and $\approx$25 respectively). The thin winner–strip above the lollipop shows that leadership still switches between budgets week-to-week rather than being monopolized by one budget, but those weekly wins translate unevenly into season-end totals — the baseline configuration converts transient advantages into the largest aggregate return. In short, ARIMA(0,0,1) with rolling re-optimization produces the strongest final outcome for the baseline setup, while tightening the budget (£60) materially reduces end-of-season payoff; intermediate budgets (£70/£80) perform between these extremes.
Figure 25: Best-performing models across families (cumulative points).
Figure 26: Family/method comparison against the benchmark ARIMA (1,0,0) Rolling (Budget = £70), shown as weekly uplift ($\Delta=$ method-benchmark) over GW 27–38. Each violin summarises the distribution of weekly $\Delta$ values, with points for individual gameweeks. Asterisks on the right margin indicate the significance of a paired Wilcoxon signed-rank test versus zero uplift (BH–FDR adjusted).
Figure 27: Absolute weekly score distributions for the top-10 methods, selected by median weekly uplift relative to the benchmark ARIMA (1,0,0) Rolling (Budget = £70). Each violin shows the distribution of weekly fantasy points over GW 27–38, with dots marking individual weeks.
Figure 28: Comparing results with (santoro2025FPL).
Figure 25 reports the best-performing method within each model family over Gameweeks 27–38. The curves show that the ARIMA(1,0,0) rolling forecaster with a £70m starting–XI budget achieves the highest cumulative score, followed closely by the weighted-average model, while the Monte Carlo and hybrid simple-average approaches trail behind. Table LABEL:tab:summary complements this figure by listing the final cumulative scores for all models shown and their comparative rankings.
To further analyze the stability of these results, Figure 26 compares each modeling family to the benchmark ARIMA (1,0,0) Rolling (£70) in terms of weekly points uplift over GW 27–38. This specific ARIMA configuration is adopted as the baseline because, as established in our cumulative performance analysis, it yields the most consistent out-of-sample results. The vertical dashed line at 0 marks parity with the benchmark; violins lying to the right correspond to methods that typically outscore it, while those centered near 0 behave similarly. ICT-based optimization and the linear-regression family deliver clear positive gains, identifying them as strong alternatives for weekly upside. In contrast, Exponential smoothing and its hybrid variant (Hybrid ExpSm 1:2) show the largest negative uplifts, with almost all weekly differences falling below zero. Monte Carlo, non-rolling ARIMA variants, and the weighted-average baselines cluster tightly around zero (or slightly negative), indicating performance comparable to, but generally outpaced by the benchmark once multiple testing is controlled for.
Figure 27 zooms in on the ten best strategies identified in the uplift analysis (Figure 26) and presents their absolute weekly scores rather than differences from the benchmark. Each horizontal violin summarizes the distribution of points across GW 27–38, with dots marking individual gameweeks. The top-ranked Hybrid ARIMA 1:2 exhibits the right-most distribution, characterized by a high median and few low-scoring weeks, indicating consistently strong performance. The various rolling ARIMA (0,0,1) and ARIMA (1,0,0) configurations cluster closely together with very similar medians and spreads, suggesting that minor budget tweaks mainly shift scores slightly without altering variability. Simple Avg Rolling also attains competitive medians but exhibits a somewhat wider violin, reflecting greater week-to-week volatility. Overall, the figure shows that the leading approaches differ by modest right-shifts in their entire score distributions rather than by occasional extreme gameweeks.
For external benchmarking, we concentrate on two strong yet relatively simple models—ARIMA(1,0,0) with a rolling window and a recency-weighted average forecaster—because they offer a good trade-off between performance and interpretability, while avoiding the extra tuning overhead of hybrid or ICT-based objectives.
Figure 28 then benchmarks these two models against the four approaches proposed in santoro2025FPL. The Greedy algorithm constructs a 15-player squad by ranking a cost–performance (“goodness”) index based on predicted mean season points, with the weekly XI chosen at random. Mean-MIP-random uses Mixed Integer Programming (MIP) on predicted mean scores to select the initial 15, again with a random weekly XI; Mean-MIP-rank keeps the MIP-optimized squad but chooses the XI each Gameweek via MIP using adjusted, GW-specific scores (form, home/away status, opponent difficulty); and Mean-MIP-online updates predictions every Gameweek and re-optimizes the XI via MIP while holding the initial squad fixed. As shown in the upper panel, our ARIMA(1,0,0)-rolling and weighted-average models outperform these baselines for most of the season. The lower panel plots the week-by-week score difference between our best model and the strongest competing method; the negative dip in Gameweek 29 is driven by data availability rather than model behavior. The only Gameweeks in which the Mean-MIP-online approach is ahead are 29 and 38; otherwise, our two models maintain the lead throughout.
Table 1: Final Cumulative Points and Ranks by Method
Method Final Total Rank
ARIMA (1,0,0) Rolling (Budget = 70) 704 1
Weighted Average 635 2
Hybrid Simple Avg. 1:2 (Higher Total Points) 561 3
Monte Carlo Simulation 545 4
(a) Weighted Average Team for Gameweek 27.
Name Team Position Value
Jordan Pickford Everton GK 4.6
Gabriel dos Santos Magalhães Arsenal DEF 5.2
William Saliba Arsenal DEF 5.7
Virgil van Dijk Liverpool DEF 6.4
Phil Foden Man City MID 8.1
Douglas Luiz Soares de Paulo Aston Villa MID 5.5
Bukayo Saka (c) Arsenal MID 9.1
Cole Palmer Chelsea MID 5.8
Mohamed Salah Liverpool MID 13.0
Ollie Watkins Aston Villa FWD 8.8
Dominic Solanke Bournemouth FWD 7.0
Matheus Cunha Wolves FWD 5.6
Jarrad Branthwaite Everton DEF 4.2
Ederson Santana de Moraes Man City GK 5.5
Conor Bradley Liverpool DEF 4.1
(b) ARIMA (1,0,0) Rolling (Budget = 70) Team for Gameweek 27.
Name Team Position Value
Đorđe Petrović Chelsea GK 4.5
Gabriel dos Santos Magalhães Arsenal DEF 5.2
William Saliba Arsenal DEF 5.7
Chris Richards Crystal Palace DEF 3.9
Phil Foden Man City MID 8.1
Douglas Luiz Soares de Paulo Aston Villa MID 5.5
Anthony Gordon Newcastle MID 6.1
Bukayo Saka (c) Arsenal MID 9.1
Cole Palmer Chelsea MID 5.8
Ollie Watkins Aston Villa FWD 8.8
Dominic Solanke Bournemouth FWD 7.0
Erling Haaland Man City FWD 14.4
Jordan Pickford Everton GK 4.6
Pedro Porro Spurs DEF 5.8
Emerson Palmieri dos Santos West Ham DEF 4.4
(a) Monte Carlo Team for Gameweek 27.
Name Team Position Value
Đorđe Petrović Chelsea GK 4.5
Gabriel dos Santos Magalhães Arsenal DEF 5.2
William Saliba Arsenal DEF 5.7
Kieran Trippier Newcastle DEF 6.9
Phil Foden Man City MID 8.1
Douglas Luiz Soares de Paulo Aston Villa MID 5.5
Son Heung-min Spurs MID 9.6
Bukayo Saka (c) Arsenal MID 9.1
Mohamed Salah Liverpool MID 13.0
Ollie Watkins Aston Villa FWD 8.8
Dominic Solanke Bournemouth FWD 7.0
Alphonse Areola West Ham GK 4.2
Conor Bradley Liverpool DEF 4.1
Cameron Archer Sheffield Utd FWD 4.3
Amari’i Bell Luton DEF 3.9
(b) Hybrid Simple Average (1:2) Team for Gameweek 27.
Name Team Position Value
Alphonse Areola West Ham GK 4.2
William Saliba Arsenal DEF 5.7
Trent Alexander-Arnold Liverpool DEF 8.6
Kieran Trippier Newcastle DEF 6.9
Phil Foden Man City MID 8.1
Douglas Luiz Soares de Paulo Aston Villa MID 5.5
Pascal Groß Brighton MID 6.5
Jarrod Bowen West Ham MID 7.7
Cole Palmer Chelsea MID 5.8
Erling Haaland (c) Man City FWD 14.4
Ollie Watkins Aston Villa FWD 8.8
Jarrad Branthwaite Everton DEF 4.2
Conor Bradley Liverpool DEF 4.1
Cameron Archer Sheffield Utd FWD 4.3
Matt Turner Nott’m Forest GK 3.9
(a) Linear Regression Team for Gameweek 27.
Name Team Position Value
Ivo Grbić Sheffield Utd GK 4.5
Gabriel dos Santos Magalhães Arsenal DEF 5.2
Daniel Muñoz (c) Crystal Palace DEF 4.5
Conor Bradley Liverpool DEF 4.1
Phil Foden Man City MID 8.1
Bukayo Saka Arsenal MID 9.1
Cole Palmer Chelsea MID 5.8
Richarlison de Andrade Spurs MID 7.2
Ollie Watkins Aston Villa FWD 8.8
Rasmus Højlund Man Utd FWD 7.2
Jayden Danns Liverpool FWD 4.5
Pablo Sarabia Wolves MID 4.7
Jarrad Branthwaite Everton DEF 4.2
Chris Richards Crystal Palace DEF 3.9
Caoimhin Kelleher Liverpool GK 3.7
(b) ICT Score Team for Gameweek 27.
Name Team Position Value
James Trafford Burnley GK 4.5
Pedro Porro Spurs DEF 5.8
Kieran Trippier Newcastle DEF 6.9
Alfie Doughty Luton DEF 4.6
Phil Foden Man City MID 8.1
Pascal Groß Brighton MID 6.5
Bruno Borges Fernandes Man Utd MID 8.2
Bukayo Saka (c) Arsenal MID 9.1
Martin Ødegaard Arsenal MID 8.4
Ollie Watkins Aston Villa FWD 8.8
Dominic Solanke Bournemouth FWD 7.0
Alphonse Areola West Ham GK 4.2
Conor Bradley Liverpool DEF 4.1
Cameron Archer Sheffield Utd FWD 4.3
Amari’i Bell Luton DEF 3.9
Tables 2(a)4(b) list the Gameweek 27 teams produced by our forecasting models (weighted average, ARIMA, Monte Carlo, and hybrid simple average) together with two additional baselines (linear regression and ICT score), revealing structural regularities in the resulting squads. Midfield-centric formations, particularly 3–5–2, appear in five of the six approaches, indicating a shared preference for midfielders’ attacking output and bonus potential under the FPL scoring system. The linear regression model is the only exception, leaning toward a 3–4–3 configuration.
Regarding captaincy, Bukayo Saka is selected as captain in four of the six models, underlining his central role in the data-driven selection process during that season. Cross-model consistency is also evident in the frequent inclusion of Phil Foden, Cole Palmer, and Ollie Watkins—the latter appearing in every best XI, in line with the Premier League’s Gameweek 27 review (“GW27 stats: Brave FPL managers rewarded for captaining Watkins”). [^7] https://www.premierleague.com/news/3917209 Arsenal assets, especially defenders Gabriel dos Santos Magalhães and William Saliba, recur prominently, reflecting a combination of strong defensive metrics and attacking involvement.
In terms of premium allocation, Mohamed Salah and Erling Haaland are each included in several line-ups but do not appear together in any of the selected squads, as their combined cost would require substantial sacrifices elsewhere unless projected returns clearly justify the investment. The goalkeeper position exhibits the greatest variability: the ARIMA and Monte Carlo models favour Đorđe Petrović, the weighted-average model opts for Jordan Pickford, the hybrid model selects Alphonse Areola, while the linear-regression and ICT-score teams choose Ivo Grbić and James Trafford, respectively. Bench selections are nonetheless broadly consistent, with Conor Bradley and Cameron Archer repeatedly appearing as low-cost enablers. Overall, the outputs converge on assets from leading clubs (notably Arsenal and Liverpool) and Aston Villa, complemented by strategic differentials in defence and goalkeeping to balance expected value with formation flexibility.
9 Conclusion
This paper introduced an integer-programming framework, together with a robust variant, for turning player-level forecasts into valid FPL lineups. The optimizer selects the starting XI, bench, and captain under budget, formation, and club constraints, using data-derived expected points as inputs. To generate these inputs, we compared averaging methods, lightweight time-series forecasting, simulation techniques, and a hybrid approach that blends recent outcomes with feature-based predictions, and we examined the impact of tighter budgets and rolling windows.
Across our experiments, recency-weighted averages and low-order ARIMA models provided strong and stable baselines. Hybrid variants often improved performance, while robust variants helped for ICT-based objectives but did not consistently outperform their non-robust counterparts. When benchmarked against the four strategies of Santorosantoro2025FPL, our ARIMA(1,0,0) rolling and weighted-average models delivered higher cumulative scores over Gameweeks 27–38, indicating that relatively simple forecasts can outperform more elaborate pipelines once embedded in an appropriate integer-programming formulation. The optimizer frequently favored a 3–5–2 shape, reflecting midfielders’ relatively high marginal contribution under the FPL scoring rules. Overall, translating noisy forecasts into constraint-aware squads produced transparent, reproducible, and practically useful improvements over benchmark strategies.
Future work includes embedding transfers and chips in a rolling-horizon setting, allowing week-to-week captaincy decisions, refining goalkeeper objectives, and enriching the feature set with external match-level data. Potential changes in FPL rules can also be incorporated directly into the optimization model.
Declaration of competing interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Author contribution
Danial Ramezani: conceptualization, data curation, methodology, experiment, visualization, writing, editing; Tai Dinh: experiment, visualization, writing, editing, validation.
Data Availability Statement
The datasets and source code are available at the link provided in the paper.
Funding Information
Not Applicable
Research Involving Human and /or Animals
Not Applicable
Informed Consent
Not Applicable